The joint probability mass function of two discrete random variables A and B is

Sca²b, a=-2,2 and b = 1,2 PA,(a,b) = 0, otherwise

(i) Are A and B are uncorrelated? (ii) Are A and B independent?
Answer #1

b=1,2 § E Pabla, b) = 1 -> ( [(-23+1 +6233x2+(2344+ (27*2] = { = c( 24)=1 c=1/24 Paras - E Pre(a,b); E žydba o t (1+2): Pigib